Chapeau Company, a U.S. corporation, operates through a branch in Champagnia. The source rules used by Champagnia are identical to those used by the United States. For 2018, Chapeau has $8,800 of gross income: $5,280 from U.S. sources and $3,520 from sources within Champagnia. The $5,280 of U.S. source income and $3,080 of the foreign source income are attributable to manufacturing activities in Champagnia (foreign branch income). The remaining $440 of foreign source income is passive category interest income. Chapeau had $2,200 of expenses other than taxes, all of which are allocated directly to manufacturing income ($880 of which is apportioned to foreign sources). Chapeau paid $422 of income taxes to Champagnia on its manufacturing income. The interest income was subject to a 10 percent withholding tax of $44.

Compute Chapeau’s total allowable foreign tax credit in 2018
Answer #1

Solution:

U.S source income Foreign income Interest income
Income 5,820 3,080 440
Less : Expenses 1,320(2,200 - 880) 770 (880 * 3,080/3,520) 110 (880 * 440/3,520 )
Income chargeable to tax 4,500 2,310 330

Tax rate = 21%(Current year U.S tax rate )

Total income = 4,500 + 2,310 + 330 = 7,140

Tax = 7,140 * 21% = 1,500

Calculation of foreign tax credit

Step 1: Foreign taxes paid = 422 + 44 = 466

Step 2:Tax to be paid = 7,140 * 21% =1,500

Step 3:Tax credit limit = Total tax payable * Foreign income / Total income

                                       = 2,310 * 2,640 / 7,140 = 855

Step 4: Step 1 or Step 3 whichever is lower

Allowable foreign tax credit = 466

Therefore allowable tax credit = 466